# Automatic Link Selection in Multi-Channel Multiple Access with Link Failures

### v1

@@ I-D Notation
[…]
I-E Definitions
In this subsection, we define some quantities that will be useful throughout. Define $s=\max\{n,m\}$. Also, define the sets
$\Delta_{l,\varepsilon}=\Big{\{}\boldsymbol{p}\in\mathbb{R}^{k}:\sum_{i=1}^{l}p_{i}=1,p_{i}\geq\varepsilon\ \forall i\in[l]\Big{\}},\text{ where }l\in\mathbb{N},$
@@ I-E Definitions
We also denote $\Delta_{l}=\Delta_{l,0}$, $\mathcal{S}^{\text{row}}=\mathcal{S}^{\text{row}}_{0}$, $\mathcal{S}^{\text{col}}=\mathcal{S}^{\text{col}}_{0}$, $\mathcal{S}^{\text{doub}}=\mathcal{S}^{\text{doub}}_{0}$, and $\mathcal{S}=\mathcal{S}_{0}$. We hide the dependence on $s$ in the notation for sets for clarity.
@@ I-F Assumptions
Before moving on to the problem, we state our main assumptions.
@@
    The function $\phi$ is concave, entrywise nondecreasing, and has bounded subgradients in $[0,1]^{n}$, i.e, $|\phi^{{}^{\prime}}_{i}(\boldsymbol{x})|\leq B\ \forall i\in[n]$, and $\boldsymbol{x}\in[0,1]^{n}$. Hence, $\phi$ is $\sqrt{n}B$-Lipschitz continuous. Also, let $\phi^{\text{max}}=\max_{\boldsymbol{x}\in[0,1]^{n}}|\phi(\boldsymbol{x})|$.
@@
    Note: This assumption is valid for most separable functions $\phi$. For instance, when $\phi$ is a proportionally fair utility function type of the form $\phi(\boldsymbol{x})=\sum_{i=1}^{n}\log(1+\beta x_{i})$, where $\beta\in\mathbb{R}_{+}$, the problem has an explicit solution.
@@ II Problem Setup
$\ \ \ \ \ \ \ (i,j)\in[n]\times[m],\tau<t,Y_{i,j}(\tau)\neq 1$ (3)
$\ \ \ \ \ \boldsymbol{Y}(t)\in\{0,1\}^{n\times m}\ \forall t\in\mathbb{N}$ (4)
$\ \ \ \ \ \sum_{i=1}^{n}Y_{i,j}(t)\leq 1\ \forall t\in\mathbb{N},j\in[m]$ (5)
$\ \ \ \ \ \sum_{j=1}^{m}Y_{i,j}(t)\leq 1\ \forall t\in\mathbb{N},i\in[n]$ (6)
@@
where constraint (2) ensures transmission decisions do not know success/failures before they happen; (3) ensures we cannot use information that is never observed. Define $s=\max\{n,m\}$ and $\phi^{\text{opt}}$ as the optimal objective value of (P1).
Lemma 1
Consider the following problem.
@@
and let $\phi^{*}$ denote the optimal objective value of the above problem. Then we have $\phi^{*}=\phi^{\text{opt}}$.
@@
We first prove that $\phi^{*}\geq\phi^{\text{opt}}$. Fix $\varepsilon>0$. Then there is a positive integer $T$ and decisions $\boldsymbol{Y}^{*}(1),\boldsymbol{Y}^{*}(2),\dots,\boldsymbol{Y}^{*}(T)$ that respect constraints (2)-(6), such that
@@
$\tilde{P}^{*}_{i,j}=\begin{cases}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{Y^{*}_{i,j}(t)\}&\text{ if }(i,j)\in[n]\times[m]\\ 0&\text{ otherwise. }\end{cases}$
Let us define $\boldsymbol{P}^{*}\in\mathcal{S}^{\text{doub}}$ such that $\boldsymbol{P}^{*}\geq\boldsymbol{\tilde{P}}^{*}$. It can be observed that such a $\boldsymbol{P}^{*}$ exists from the definition of $\boldsymbol{\tilde{P}}^{*}$ and the constraints (4)-(6). Let $\boldsymbol{\gamma}^{*}\in[0,1]^{n}$ such that $\gamma^{*}_{i}=\sum_{j=1}^{m}q_{i,j}P^{*}_{i,j}$ for $i\in[n]$. It is easy see that $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ satisfy constraints (9)-(11). Also, we have
@@
where (a) follows since $\phi^{*}$ is the optimal objective value of (P2) and $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ is feasible for (P2), (b) follows from (13) and the entrywise nondecreasing property of $\phi$ (see assumption A1) and (c) follows from (12). The above is true for all $\varepsilon>0$. Hence, we have $\phi^{*}\geq\phi^{\text{opt}}$ as desired.
Now, we prove that $\phi^{*}\leq\phi^{\text{opt}}$. Let $\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*}$ denote the optimal solution for (P2). Using Birkhoff-von Neumann Theorem [50], notice that there exists $r\in\mathbb{N}$ and permutation matrices $\boldsymbol{M}^{1},\dots,\boldsymbol{M}^{r}$ such that $\boldsymbol{P}^{*}=\sum_{l=1}^{r}s_{l}\boldsymbol{M}^{l}$, and $\boldsymbol{s}\in\Delta_{r}$. In each time slot, we simply sample $l_{t}\sim\boldsymbol{s}(t)$ and set $\boldsymbol{Y}(t)$ according to $Y_{i,j}(t)=M^{l_{t}}_{i,j}$ for $i\in[n],j\in[m]$. It can be easily seen that this policy yields an objective value of $\phi^{*}$ for (P1). Hence, we are done. ∎
III Adaptive Algorithm
Now, we develop our first algorithm. The idea is to find $\boldsymbol{\tilde{P}}(t)$ as a column stochastic matrix and a row stochastic matrix alternatively in odd and even slots. Then we obtain $\boldsymbol{P}(t)$ by approximating $\boldsymbol{\tilde{P}}(t)$ by a doubly stochastic matrix, using the rounding trick similar to the one introduced in [33], after which we sample a permutation matrix from $\boldsymbol{P}(t)$ using Birkhoff-von Neumann Decomposition [50]. To do the assignment in time slot $t-1$, we discard the last $s-n$ rows or last $s-m$ columns of the permutation matrix, depending on whether $m$ or $n$ is larger.
@@ III Adaptive Algorithm
1 Initialize $\boldsymbol{\tilde{P}}(1)\in\mathcal{S}^{\text{doub}}_{\varepsilon}$ and the virtual queues $\boldsymbol{Q}(1)\in[0,BV+1]^{n}$ arbitrarily.
@@
3       Set $\boldsymbol{P}(t)=\text{ROUND}(\boldsymbol{\tilde{P}}(t))$ (This function yields $\boldsymbol{P}(t)\in\mathcal{S}^{\text{doub}}_{\varepsilon/s}$ ).
4
5      Using Birkhoff-von Neumann Decomposition [50], find $r\in\mathbb{N}$ and permutation matrices $\boldsymbol{M}^{1},\dots,\boldsymbol{M}^{r}$ such that $\boldsymbol{P}(t)=\sum_{l=1}^{r}s_{l}(t)\boldsymbol{M}^{l}$, and $\boldsymbol{s}(t)\in\Delta_{r}$.
6       Sample $l_{t}\sim\boldsymbol{s}(t)$ and take action $\boldsymbol{Y}(t)$, where $Y_{i,j}(t)=M^{l_{t}}_{i,j}$ for all $i\in[n]$, $j\in[m]$, and receive $\boldsymbol{S}(t)\odot\boldsymbol{Y}(t)$ as feedback.
8       Find $\boldsymbol{\gamma}(t+1)\in[0,1]^{n}$, and $\boldsymbol{\tilde{P}}(t+1)\in\mathcal{Q}$ using
$\boldsymbol{\tilde{P}}(t+1)=\arg\min_{\boldsymbol{P}\in\mathcal{Q}}\Big{[}-\sum_{i=1}^{n}\sum_{j=1}^{m}Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad+\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{\tilde{P}}(t))\Big{]},$ (14)
where $\mathcal{Q}=\mathcal{S}^{\text{col}}_{\varepsilon}$ if $t$ is even, and $\mathcal{Q}=\mathcal{S}^{\text{row}}_{\varepsilon}$ if $t$ is odd, and the divergence $D$ is defined by
$D(\boldsymbol{X}\lVert\boldsymbol{Y})=\sum_{i=1}^{s}\sum_{j=1}^{s}X_{i,j}\log\left(\frac{X_{i,j}}{Y_{i,j}}\right).$ (15)
for all $\boldsymbol{X},\boldsymbol{Y}\in\mathcal{S}.$

### v2

@@ I-D Notation
For integers $a,b$, we use $[a:b]$ to denote the set of integers between $a,b$ inclusive. We use $[a]=[1:a]$. For $\boldsymbol{a}\in\mathbb{R}^{k}$, $\lVert\boldsymbol{a}\rVert=\sqrt{\sum_{i=1}^{k}a_{i}^{2}}$, $\lVert\boldsymbol{a}\rVert_{1}=\sum_{i=1}^{k}|a_{i}|$, and $\boldsymbol{a}]_{+}\in\mathbb{R}^{k}$ is the vector with $i$-th entry $\max\{a_{i},0\}$. We use $\boldsymbol{1}_{k}$ to denote the $k$-dimensional vector of ones. When the dimension is clear from context, we use $\boldsymbol{1}$ instead of $\boldsymbol{1}_{k}$. For vectors $\boldsymbol{a,b}\in\mathbb{R}^{k}$, $\boldsymbol{c}=\boldsymbol{a}\odot\boldsymbol{b}\in\mathbb{R}^{k}$ is defined such that $c_{i}=a_{i}b_{i}$ for all $i\in[k]$. For matrices $\boldsymbol{A,B}\in\mathbb{R}^{k\times l}$ we use $\lVert\boldsymbol{A}\rVert=\sqrt{\sum_{i=1}^{k}\sum_{j=1}^{l}A_{i,j}^{2}}$, $\lVert\boldsymbol{A}\rVert_{1}=\sum_{i=1}^{k}\sum_{j=1}^{l}|A_{i,j}|$, and $\boldsymbol{C}=\boldsymbol{A}\odot\boldsymbol{B}\in\mathbb{R}^{k\times l}$ is defined such that $C_{i,j}=A_{i,j}B_{i,j}$ for all $i\in[k]$ and $j\in[l]$.
I-E Definitions and Preliminaries
In this subsection, we define some quantities that will be useful throughout. Define $s=\max\{n,m\}$. Also,
$\Delta^{l}_{\varepsilon}=\left\{\boldsymbol{p}\in\mathbb{R}^{l}:\sum_{i=1}^{l}p_{i}=1,p_{i}\geq\varepsilon\ \forall i\in[l]\right\},\text{ where }l\in\mathbb{N},$
@@ I-E Definitions and Preliminaries
We also denote $\Delta^{l}=\Delta^{l}_{0}$, $\mathcal{S}^{\text{row}}=\mathcal{S}^{\text{row}}_{0}$, $\mathcal{S}^{\text{col}}=\mathcal{S}^{\text{col}}_{0}$, $\mathcal{S}^{\text{doub}}=\mathcal{S}^{\text{doub}}_{0}$, and $\mathcal{S}=\mathcal{S}_{0}$. Also, let $\Delta^{l}_{+}=\Delta^{l}\cap(0,1]^{l}$, $\mathcal{S}^{\text{doub}}_{+}=\mathcal{S}^{\text{doub}}\cap(0,1]^{s\times s}$ and $\mathcal{S}_{+}=\mathcal{S}\cap(0,1]^{s\times s}$. We hide the dependence on $s$ in the notation for sets for clarity.
$\beta$-approximation
Given $\beta>0$, and matrix $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{+}$, we say matrix $\hat{P}\in\mathcal{S}^{\text{doub}}_{+}$ is a $\beta$-approximation of $\boldsymbol{P}$ if
$e^{-\beta}\leq\frac{\hat{P}_{j,j}}{P_{i,j}}$ (2)
hold for all $i,j\in[s]$.
KL-divergence
For $\boldsymbol{p}\in\Delta^{s}$ and $\boldsymbol{q}\in\Delta^{s}_{+}$, we define the Kullback–Leibler (KL) divergence as
$D_{\mathrm{KL}}(\boldsymbol{p}\lVert\boldsymbol{q})=\sum_{i=1}^{s}p_{i}\ln\left(\frac{p_{i}}{q_{i}}\right),$ (3)
with the convention that $0\ln(0/q)=0$ for any $q>0$. For matrices $\boldsymbol{P}\in\mathcal{S}$ and $\boldsymbol{Q}\in\mathcal{S}_{+}$, we define the (entrywise) matrix KL divergence by
$D(\boldsymbol{P}\lVert\boldsymbol{Q})=\sum_{i=1}^{s}\sum_{j=1}^{s}P_{i,j}\ln\left(\frac{P_{i,j}}{Q_{i,j}}\right),$ (4)
where again $0\ln(0/q)=0$ for $q>0$.
Sampling from $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}$
We define the process of sampling a permutation matrix $\boldsymbol{Y}$ from a doubly stochastic matrix $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}$. By the Birkhoff–von Neumann theorem [12], there exists an integer $r\in\mathbb{N}$ and permutation matrices $\boldsymbol{M}^{1},\dots,\boldsymbol{M}^{r}$ such that $\boldsymbol{P}=\sum_{l=1}^{r}s_{l}\boldsymbol{M}^{l}$, where $\boldsymbol{s}\in\Delta^{r}$ is a probability vector over $[1:r]$. Consider the following sampling procedure: draw an index $a\in[1:r]$ according to the distribution $\boldsymbol{s}$ and set $\boldsymbol{Y}=\boldsymbol{M}^{a}$. Then, $\mathbb{E}\!\left\{\boldsymbol{Y}\mid\boldsymbol{P}\right\}=\boldsymbol{P}$.
@@ I-F Assumptions
We state our main assumptions.
@@
    The function $\phi$ is concave, entrywise nondecreasing, and has bounded subgradients in $0,1]^{n}$, i.e, $|\phi^{{}^{\prime}}_{i}(\boldsymbol{x})|\leq B\ \forall i\in[n]$, and $\boldsymbol{x}\in[0,1]^{n}$. Hence, $\phi$ is $\sqrt{n}B$-Lipschitz continuous. Also, let $\phi^{\text{max}}=\max_{\boldsymbol{x}\in[0,1]^{n}}|\phi(\boldsymbol{x})|$.
@@
    Note: This is valid for most separable functions $\phi$. For instance, when $\phi$ is a proportionally fair utility function type of the form $\phi(\boldsymbol{x})=\sum_{i=1}^{n}\log(1+\beta x_{i})$, where $\beta\in\mathbb{R}_{+}$, the problem in A2 has an explicit solution.
@@ II Problem Setup
$\text{s.t. }\boldsymbol{Y}(t),\text{ and }\boldsymbol{S}(\tau)\text{ are independent for all }$
@@
$\ \ \ \ \ \ \ (i,j)\in[s]\times[s],\tau<t,Y_{i,j}(\tau)\neq 1$ (7)
$\ \ \ \ \ \boldsymbol{Y}(t)\in\{0,1\}^{s\times s}\ \forall t\in\mathbb{N}$ (8)
$\ \ \ \ \ \sum_{i=1}^{s}Y_{i,j}(t)\leq 1\ \forall t\in\mathbb{N},j\in[s]$ (9)
$\ \ \ \ \ \sum_{j=1}^{s}Y_{i,j}(t)\leq 1\ \forall t\in\mathbb{N},i\in[s]$ (10)
@@
where constraint (6) ensures transmission decisions do not know success/failures before they happen; (7) ensures we cannot use information that is never observed. Define $\phi^{\text{opt}}$ as the optimal objective value of (P1).
II-A Deterministic Problem
Consider the deterministic problem given by
@@ II-A Deterministic Problem
Let $\phi^{*}$ denote the optimal objective value of the above problem. In Lemma 1, we show that $\phi^{*}\geq\phi^{\text{opt}}$. This lemma will be useful in establishing the performance guarantees of our algorithms. Let us define $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ as the optimal solution Problem (P2).
Note: More generally, it can be shown that $\phi^{*}=\phi^{\text{opt}}$. If $q_{i,j}$ were known, Problem (P2) could be solved to obtain $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$. The stationary policy that samples a permutation matrix from $\boldsymbol{P}^{*}$ in each time slot (see Section I-E-c) and uses it for assignment achieves an objective value of $\phi^{*}$ for Problem (P1). However, even when $q_{i,j}$ is unknown, Section IV establishes that our algorithm attains an objective value arbitrarily close to $\phi^{*}$, thereby implying that $\phi^{\text{opt}}\geq\phi^{*}$.
Lemma 1
We have $\phi^{*}\geq\phi^{\text{opt}}$, where $\phi^{\text{opt}}$ and $\phi^{*}$ are the optimal objective values of Problems (P1) and  (P2), respectively.
@@
Fix $\varepsilon>0$. Then there is a positive integer $T$ and decisions $\boldsymbol{Y}^{*}(1),\boldsymbol{Y}^{*}(2),\dots,\boldsymbol{Y}^{*}(T)$ that respect constraints (6)-(10), such that
@@
$\tilde{P}^{*}_{i,j}=\begin{cases}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{Y^{*}_{i,j}(t)\}&\text{ if $(i,j)\in[n]\times[m]$}\\ 0&\text{ otherwise}\end{cases}$ (17)
Constraints (8)–(10) ensure the $s\times s$ matrix $\boldsymbol{\tilde{P}}^{*}$ has nonnegative entries with all row and column sums less than or equal to 1. A well known result of Von-Neumann ensures that there is a matrix $\boldsymbol{P}^{*}\in\mathcal{S}^{\text{doub}}$ such that $\tilde{P}^{*}_{i,j}\leq P^{*}_{i,j}$ for all $i,j\in[s]$. Let $\boldsymbol{\gamma}^{*}\in[0,1]^{n}$ such that $\gamma^{*}_{i}=\sum_{j=1}^{m}q_{i,j}P^{*}_{i,j}$ for $i\in[n]$. It is easy see that $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ satisfy constraints (13)-(15). Also, we have
@@
where (a) follows since $\phi^{*}$ is the optimal objective value of (P2) and $(\boldsymbol{P}^{*},\boldsymbol{\gamma}^{*})$ is feasible for (P2), (b) follows from (18) and the entrywise nondecreasing property of $\phi$ (see assumption A1) and (c) follows from (16). The above is true for all $\varepsilon>0$. Hence, we have $\phi^{*}\geq\phi^{\text{opt}}$ as desired. ∎
III Intuition Behind Algorithms
Before moving onto the algorithms, we describe the general intuition unifying all the algorithms. Consider the partial Lagrangian of problem (P2) defined as
$f(\boldsymbol{Q},\boldsymbol{P},\boldsymbol{\gamma})=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}\!\left(\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right),$ (19)
obtained by dualizing constraint (15), where $Q_{i}\geq 0$ are the Lagrange multipliers associated with (15), and $V>0$ is a tunable parameter that scales the utility term without affecting the optimal solution of (P2). If the success probabilities $q_{i,j}$ were known, then for a suitably chosen parameter $V$, the iterative primal update
$(\boldsymbol{\gamma}(t+1),\boldsymbol{P}(t))$ $=\arg\min_{\begin{subarray}{c}\boldsymbol{P}\in\mathcal{S}^{\text{doub}}\\ \boldsymbol{\gamma}\in[0,1]^{n}\end{subarray}}f(\boldsymbol{Q}(t),\boldsymbol{P},\boldsymbol{\gamma})$ (20)
$=\arg\min_{\boldsymbol{P},\boldsymbol{\gamma}}\left\{-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\left[\gamma_{i}-\sum_{j=1}^{m}q_{i,j}P_{i,j}\right]\right\},$
together with the dual update
$Q_{i}(t+1)=\left[Q_{i}(t)+\gamma_{i}(t+1)-\sum_{j=1}^{m}q_{i,j}P_{i,j}(t)\right]_{+},$ (21)
for all $i\in[n]$, guarantees
$\phi(\bar{\boldsymbol{\gamma}}(T))$ $\geq\phi^{\text{opt}}-\frac{o(T)}{T},$ (22)
and
$\bar{\gamma}_{i}(T)$ $\leq\sum_{j=1}^{m}q_{i,j}\bar{P}_{i,j}(T)+\frac{o(T)}{T},\quad\forall i\in[n],$ (23)
where the time averages are defined as $(\bar{\boldsymbol{\gamma}}(T),\bar{\boldsymbol{P}}(T))=\frac{1}{T}\sum_{t=1}^{T}(\boldsymbol{\gamma}(t),\boldsymbol{P}(t))$. The result follows from standard primal–dual convergence arguments [39]. Hence, we can get arbitrarily close to the optimal solution of (P2).
We can either find $\boldsymbol{\bar{P}}(T)$ offline using the above approach, and implement the fixed policy of sampling an assignment from $\boldsymbol{\bar{P}}(T)$ (see Section I-E-c) in each time slot, or we can implement an online procedure where we sample an assignment from $\boldsymbol{P}(t)$ in time slot $t$. Both of these approaches will guarantee (1) given that $q_{i,j}$ values are known.
All of our algorithms follow the structure described above. However, in our scenario, we do not know the values of $q_{i,j}$. Hence, they have to be estimated using the bandit feedback. For this task, in the adaptive algorithms, we use the EXP3.S importance sampling based estimation, whereas in the non-adaptive algorithm, we use UCB-based estimation. For each $i\in[n]$, the sequence $Q_{i}(1),Q_{i}(2),\dots$ is referred to as a virtual queue, since the dual update (21) resembles a standard queueing recursion. Observe that the primal update (20) decomposes into separate decisions over $\boldsymbol{\gamma}$ and $\boldsymbol{P}$. This decomposition is the motivation behind the introduction of the auxiliary variable $\boldsymbol{\gamma}$. In particular, it enables us to decouple the matching decisions from the utility optimization.
IV Adaptive MAC Algorithm
Now, we develop our first algorithm. Similar to the description of the known $q_{i,j}$ setting in Section III, the idea is to implement an online algorithm to find a doubly stochastic matrix $\boldsymbol{P}(t)$ in each time using the success/failure feedback. However, this is more challenging since we do not know $q_{i,j}$, and we only receive bandit feedback on success/failures. In addition, we require the algorithm to adapt when the success failure probabilities change. To achieve this, we combine EXP3.S based importance sampling idea with Lyapunov optimization.
In each time slot, we solve a convex optimization problem to obtain an intermediate matrix $\boldsymbol{\tilde{P}}(t)$. The problem cannot be solved exactly. Hence, we find an approximation $\boldsymbol{\hat{P}}(t)$ of $\boldsymbol{\tilde{P}}(t)$. We then find the doubly stochastic matrix $\boldsymbol{P}(t)$ by mixing $\boldsymbol{\hat{P}}(t)$ with the all one matrix. This step is vital in ensuring adaptiveness. We then sample $\boldsymbol{Y}(t)$ from $\boldsymbol{P}(t)$ using Birkhoff-von Neumann Decomposition [12]. We assume that if $i>n$ or $j>m$, $Y_{i,j}(t)$ corresponds to a fictional link. Formally, we extend the notation by setting $S_{i,j}(t)=0$ and $q_{i,j}=0$ for $i>n$ and $j>m$. In this algorithm, we use the loss-based importance sampling estimator
$\hat{S}_{i,j}(t)=1-\frac{(1-S_{i,j}(t))Y_{i,j}(t)}{P_{i,j}(t)}$ (24)
of $q_{i,j}$. This enables fast convergence guarantees.
In Algorithm 1, we provide the algorithm for the task. The algorithm uses parameters $V>0$, $\theta>0$, $\varepsilon\in(0,1/2]$, and $\eta>0$.
@@ IV Adaptive MAC Algorithm
1Initialize $\boldsymbol{P}(1)=\frac{1}{s}\boldsymbol{1}\boldsymbol{1}^{\top}\in\mathcal{S}^{\text{doub}}$ and the virtual queues $\boldsymbol{Q}(1)=\boldsymbol{0}\in\mathbb{R}^{n}$.
@@
3   Sample $\boldsymbol{Y}(t)$ from $\boldsymbol{P}(t)$ (see Section I-E-c), and receive $\boldsymbol{S}(t)\odot\boldsymbol{Y}(t)$ as feedback.
4   Compute the estimator $\boldsymbol{\hat{S}}(t)$ for $\boldsymbol{S}(t)$ using $\hat{S}_{i,j}(t)=1-\frac{(1-S_{i,j}(t))Y_{i,j}(t)}{P_{i,j}(t)}$ for all $i,j\in[s]$ where we define $S_{i,j}(t)=0$ whenever $i>n$ or $j>m$.
5   Define
$\boldsymbol{\gamma}(t+1)=\arg\min_{\boldsymbol{\gamma}\in[0,1]^{n}}\Big[-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)\gamma_{i}\Big].$ (25)
6   Find $\boldsymbol{\hat{P}}(t+1)$ as a $(\theta/s)$-approximation of $\boldsymbol{\tilde{P}}(t+1)$ where
$\boldsymbol{\tilde{P}}(t+1)$ $=\arg\min_{\boldsymbol{P}\in\mathcal{S}^{\text{doub}}}\Big[-\sum_{i=1}^{n}\sum_{j=1}^{s}Q_{i}(t)\hat{S}_{i,j}(t)P_{i,j}+\frac{1}{\eta}D(\boldsymbol{P}\lVert\boldsymbol{P}(t))\Big],$ (26)
where $\beta$-approximation is defined in (2) and the divergence $D(\cdot\lVert\cdot)$ is defined in (4).
7   Mix
$\boldsymbol{P}(t+1)=(1-\varepsilon)\boldsymbol{\hat{P}}(t+1)+\frac{\varepsilon}{s}\boldsymbol{1}\boldsymbol{1}^{\top}$ (27)
8